# Reconstructions of Einstein-Aether Gravity from Barrow Agegraphic and New Barrow Agegraphic Dark Energy models: Examinations and Observational Limits

### v1

@@ 5.2 Pantheon+SH0ES Supernova Dataset:
Table 4: Best-fit Parameters for BADE, NBADE, and $\Lambda$CDM Models
Model Parameter Pantheon+SH0ES CC+BAO
BADE $H_{0}$ $72.968_{-0.041}^{+0.023}$ $69.059_{-0.903}^{+1.540}$
$\Omega_{m0}$ $0.314_{-0.066}^{+0.034}$ $0.326_{-0.060}^{+0.040}$
$\omega_{m}$ $0.139_{-0.025}^{+0.036}$ $0.005_{-0.029}^{+0.038}$
$\Delta$ $0.324_{-0.201}^{+0.322}$ $0.389_{-0.275}^{+0.275}$
$M$ $0.771_{-0.166}^{+0.163}$ $0.714_{-0.136}^{+0.180}$
NBADE $H_{0}$ $72.992_{-0.014}^{+0.006}$ $69.508_{-1.723}^{+2.348}$
$\Omega_{m0}$ $0.309_{-0.002}^{+0.001}$ $0.272_{-0.064}^{+0.066}$
$\omega_{m}$ $0.132_{-0.014}^{+0.015}$ $-0.028_{-0.043}^{+0.042}$
$\Delta$ $0.467_{-0.286}^{+0.351}$ $0.451_{-0.243}^{+0.331}$
$M$ $0.553_{-0.243}^{+0.270}$ $0.626_{-0.301}^{+0.273}$
$\Lambda$CDM $H_{0}$ $73.079_{-0.158}^{+0.157}$ $69.989_{-1.197}^{+0.876}$
$\Omega_{m0}$ $0.341_{-0.014}^{+0.006}$ $0.268_{-0.015}^{+0.017}$
@@ 6 Results and Discussion
EoS Parameter and the $\omega^{\prime}_{DE}-\omega_{DE}$ Phase Plane: The dynamical characteristics of the dark energy equation of state (EoS) parameter $\omega_{DE}$ are presented in Fig. 10. The analysis reveals a consistently negative parametric trajectory throughout cosmic history, with an upward trend as the universe approaches the present epoch (decreasing $z$). This characteristic pattern indicates the emergence and subsequent dominance of the dark energy component. The observed behavior in both theoretical frameworks - BADE and NBADE - exhibits signatures consistent with quintessence-like dark energy dynamics. Contemporary measurements yield distinct present-day values: the BADE construction predicts $\omega_{DE0}=-0.8718$, while the NBADE formalism suggests $\omega_{DE0}=-0.7841$. These theoretical predictions demonstrate subtle variations from contemporary observational constraints [133].
Figure 10: Plot of evolution trajectory of EoS parameter for dark energy $\omega_{DE}(z)$ for the fig: (a) BADE model and fig: (b) NBADE model.
The phase plane of $\omega^{\prime}_{DE}-\omega_{DE}$ depicted in Fig. 11 reveals a freezing region for both models, indicating an enhancement in the acceleration of the universe’s expansion.
@@
Square of the Speed of Sound: The plot in Fig. 13 demonstrates that the square of the sound velocity $v_{s}^{2}$ is positive for $z>1$ and negative for $z<1$ in the case of the BADE model and NBADE model. This indicates that the BADE model and NBADE model are stable at higher redshifts but become unstable at lower redshift values. It should be emphasized that both models are classified as partially stable.
@@
Statefinder Parameter: The analysis of dynamical parameters in Figure 14 reveals the evolutionary trajectory in the $(r,s^{*})$ plane. During early cosmic epochs, both cosmological frameworks exhibit characteristics defined by $(r>1)$ and $(s^{*}<0)$, corresponding to quintessence behavior. The evolutionary path intersects the $\Lambda$CDM fixed point at $(r,s^{*})=(1,0)$, subsequently entering a domain characterized by $(r<1)$ and $(s^{*}>0)$, which is indicative of Chaplygin gas behavior. The dynamical evolution portrayed in Figure 15 through the $(r,q)$ plane demonstrates the cosmic transition from decelerated to accelerated expansion. This transformation reflects the shift from matter dominance to dark energy prevalence. Notable is the absence of convergence to either the steady-state configuration $(r=1,q=-1)$ or the SCDM point $(r=1,q=\frac{1}{2})$. Instead, the asymptotic behavior of both models tends toward the de Sitter state, characterized by $q=-1$. In the context of Einstein-Aether gravity reconstruction, the BADE framework exhibits more rapid convergence compared to its NBADE counterpart.
@@
7 Conclusion
Our investigation presents a comprehensive analysis of Einstein-Aether gravity through two distinct theoretical frameworks: the Barrow Agegraphic Dark Energy (BADE) and New Barrow Agegraphic Dark Energy (NBADE) formulations. The reconstruction methodology initiated with the derivation of characteristic functions $F(K)$ in Einstein-Aether gravity by establishing equivalence between the energy densities of BADE, NBADE, and Einstein-Aether constructions. The analysis demonstrates that $F(K)$ exhibits analytical properties under specific parametric constraints involving the Barrow exponent $\Delta$. A notable distinction emerges in the functional behavior: while the BADE framework yields a monotonically increasing $F(K)$, the NBADE construction manifests decreasing behavior with increasing $K$.
Statistical validation employed Markov Chain Monte Carlo (MCMC) techniques, integrating observational constraints from 57 CC+BAO measurements and 1701 Pantheon+SH0ES datapoints. The derived Hubble parameter estimations demonstrate strong concordance with Planck Collaboration measurements, while exhibiting minor tensions with SH0ES findings. The redshift evolution of $H(z)$ reveals remarkable agreement with empirical data, though distinctive features emerge at higher redshifts $(z>0.5)$, suggesting unique physical mechanisms operating in these cosmological epochs.
The deceleration-acceleration transition manifests at characteristic redshifts $z_{t}\approx 0.7568$ (BADE) and $z_{t}\approx 0.6561$ (NBADE), displaying consistency with combined constraints from SNIa, CMB, and LSS observations. Higher-order kinematic parameters - jerk, snap, and lerk - provide additional insights into cosmic acceleration dynamics. The dark energy equation of state $\omega_{DE}$ maintains negative values throughout cosmic evolution, indicating quintessence-like behavior in both theoretical frameworks. The $\omega^{\prime}_{DE}-\omega_{DE}$ phase space analysis reveals freezing behavior, suggesting accelerated cosmic expansion.
The normalized density parameters $\Omega(z)$ identify a matter-dark energy dominance transition near $z\approx 0.7$. Stability analysis through sound velocity $v_{s}^{2}$ indicates epoch-dependent behavior: stability at high redshifts transitioning to instability in recent cosmic history. Phase space trajectories in $(r,s^{*})$ and $(r,q)$ planes characterize the evolution from quintessence to Chaplygin gas regimes, with asymptotic approach toward de Sitter behavior. The Om diagnostic reveals dynamic dark energy characteristics, demonstrating transitions between phantom and quintessence phases.
Parameter estimation yields cosmologically viable values for matter density $\Omega_{m0}$, matter equation of state $\omega_{m}$, Barrow exponent $\Delta$, coupling constant $M$, and cosmic age $t_{\text{age}}$, demonstrating consistency with contemporary observational constraints and theoretical predictions within Einstein-Aether gravity.
While our research has provided valuable insights, several avenues for future research remain open. Further investigation into the physical implications of the Barrow exponent $\Delta$ and its relationship to cosmic curvature is warranted. The exploration of the transition between phantom-like and quintessence-like dark energy regimes observed in the Om diagnostic could yield additional insights. Development of more sophisticated models that can account for the slight deviations from $\Lambda$CDM at higher redshifts and investigation of the partial stability of the models and its implications for the long-term evolution of the universe are also important areas for future work.
An intriguing direction for future research is the introduction of a 4-parameter generalized entropy, such as the Nojiri-Odintsov-Paul entropy [135], which reduces to Barrow entropy under specific parameter limits. This approach could provide a broader theoretical framework for our current proposal. Additionally, investigating the possibility that the Barrow exponent $\Delta$ might vary with energy scale [136, 137] could offer novel implications for early universe cosmology and potentially lead to the consideration of a time-varying $\Delta$, although this diverges from the agegraphic dark energy context addressed in this work. Additionally, one can introduce an interaction term to incorporate the interaction between dark matter and the dark energy shown in [73] to further modify the model with real scenarios.
To further refine our understanding, the incorporation of more extensive cosmological datasets, including CMB, gravitational wave, and Planck data, could provide tighter constraints on model parameters and offer a more comprehensive test of the model’s validity across different cosmological epochs. Lastly, running simulations of large-scale structure (LSS) formation using this model would significantly enhance our understanding of the universe’s evolution, particularly in contrast to the standard $\Lambda$CDM paradigm, and could provide additional observational tests for the model.
In summary, our reconstruction of Einstein-Aether gravity through the BADE and NBADE models yields profound insights into the characteristics of dark energy and the history of the universe’s expansion. Although these models demonstrate encouraging consistency with existing observational data, they also raise new questions and possibilities for future investigations in cosmology and gravitational physics. The suggested future directions, spanning from theoretical extensions to enhanced observational assessments, present intriguing opportunities for further refining our grasp of the fundamental aspects of the universe and its evolution.

### v2

@@ 5.2 Pantheon+SH0ES Supernova Dataset:
Table 5: Best-fit Parameters and Prior Ranges for BADE, NBADE, and $\Lambda$CDM Models
Model Parameter Pantheon+SH0ES CC+BAO Prior Range
BADE $H_{0}$ $72.968_{-0.041}^{+0.023}$ $69.059_{-0.903}^{+1.540}$ [60, 70]6070[60,\ 70][ 60 , 70 ]
$\Omega_{m0}$ $0.314_{-0.066}^{+0.034}$ $0.326_{-0.060}^{+0.040}$ [0.1, 0.3]0.10.3[0.1,\ 0.3][ 0.1 , 0.3 ]
$\omega_{m}$ $0.139_{-0.025}^{+0.036}$ $0.005_{-0.029}^{+0.038}$ [0.5, 0.5]0.50.5[-0.5,\ 0.5][ - 0.5 , 0.5 ]
$\Delta$ $0.324_{-0.201}^{+0.322}$ $0.389_{-0.275}^{+0.275}$ [0, 1]01[0,\ 1][ 0 , 1 ]
$M$ $0.771_{-0.166}^{+0.163}$ $0.714_{-0.136}^{+0.180}$ [2, 2]22[-2,\ 2][ - 2 , 2 ]
NBADE $H_{0}$ $72.992_{-0.014}^{+0.006}$ $69.508_{-1.723}^{+2.348}$ [60, 72]6072[60,\ 72][ 60 , 72 ]
$\Omega_{m0}$ $0.309_{-0.002}^{+0.001}$ $0.272_{-0.064}^{+0.066}$ [0.1, 0.3]0.10.3[0.1,\ 0.3][ 0.1 , 0.3 ]
$\omega_{m}$ $0.132_{-0.014}^{+0.015}$ $-0.028_{-0.043}^{+0.042}$ [0.1, 1.2]0.11.2[-0.1,\ 1.2][ - 0.1 , 1.2 ]
$\Delta$ $0.467_{-0.286}^{+0.351}$ $0.451_{-0.243}^{+0.331}$ [0, 1]01[0,\ 1][ 0 , 1 ]
$M$ $0.553_{-0.243}^{+0.270}$ $0.626_{-0.301}^{+0.273}$ [0.09, 1.0]0.091.0[0.09,\ 1.0][ 0.09 , 1.0 ]
$\Lambda$CDM $H_{0}$ $73.079_{-0.158}^{+0.157}$ $69.989_{-1.197}^{+0.876}$ [60, 70]6070[60,\ 70][ 60 , 70 ]
$\Omega_{m0}$ $0.341_{-0.014}^{+0.006}$ $0.268_{-0.015}^{+0.017}$ [0.1, 0.3]0.10.3[0.1,\ 0.3][ 0.1 , 0.3 ]
@@ 6 Results and Discussion
EoS Parameter and the $\omega^{\prime}_{DE}-\omega_{DE}$ Phase Plane: The dynamical characteristics of the dark energy equation of state (EoS) parameter $\omega_{DE}$ are presented in Fig. 10. The analysis reveals a consistently negative parametric trajectory throughout cosmic history, with an upward trend as the universe approaches the present epoch (decreasing $z$). This characteristic pattern indicates the emergence and subsequent dominance of the dark energy component. The observed behavior in both theoretical frameworks - BADE and NBADE - exhibits signatures consistent with quintessence-like dark energy dynamics. Contemporary measurements yield distinct present-day values: the BADE construction predicts $\omega_{DE0}=-0.8718$, while the NBADE formalism suggests $\omega_{DE0}=-0.7841$. These theoretical predictions demonstrate close alignments from contemporary observational constraints from DES-SN5YR [142] and DESI-VI [143]. In other words, the models are in the quintessence regime today. However, both $\omega_{DE}(z)$ curves move towards $-1$ as redshift decreases, indicating an asymptotic crossing into the phantom regime ($\omega_{DE}<-1$) in the future. This crossing of the $w=-1$ divide is not abrupt but happens at $z<0$ (i.e. it lies in the extrapolated future). Such behavior is not unique to the Einstein-Aether reconstructions: it is well known, for example, that Barrow holographic dark energy can similarly evolve from quintessence to phantom as parameters change [144]. Indeed, many modified gravity or interacting dark energy scenarios can achieve an effective phantom crossing without pathologies [22]. In contrast, a simple canonical quintessence model has $\omega_{DE}>-1$ at all times, and models like the generalized Chaplygin gas smoothly approach $\omega_{DE}\to-1$ from above (never truly crossing) at late times. Thus our models’ eventual phantom-like fate (driven by the Barrow exponent effects and Aether coupling) is a novel prediction, but it parallels trends seen in other exotic DE frameworks.
Figure 10: Plot of evolution trajectory of EoS parameter for dark energy $\omega_{DE}(z)$ for the fig: (a) BADE model and fig: (b) NBADE model. (c) A precise comparison with other models.
The phase-plane trajectory in the $(\omega_{DE},\omega^{\prime}_{DE})$ space (Fig. 11) lies in the so-called freezing region: the curves move toward $\omega_{DE}=-1$ with a slowly decreasing derivative (i.e. $\omega^{\prime}_{DE}<0$). This indicates that dark energy is dynamically settling into a cosmological constant–like state as time passes. Freezing behavior is characteristic of many tracking quintessence models [145] and differs from “thawing” models where $w$ evolves away from $-1$. Thus, even though our present-day $\omega_{DE}$ is slightly above $-1$, its future trend is toward $-1$, implying ongoing acceleration that will asymptotically resemble de Sitter expansion.
@@
Square of the Speed of Sound: The models’ viability as cosmological fluids can be assessed via the squared sound speed $v_{s}^{2}$ (Fig. 13). For both EA-BADE and EA-NBADE, $v_{s}^{2}$ is positive at high redshift and becomes negative at late times, indicating that the reconstructed Aether–DE mixture is stable (no perturbation growth) in the past but develops an instability in the recent universe. In other words, both models are partially stable: small early perturbations do not grow, but late-time perturbations could. This echoes findings in some agegraphic or holographic models where stability is marginal. This behavior is consistent with recent studies on agegraphic and holographic dark energy models, which report similar partial stability, with stability at high redshifts and marginal or unstable conditions at late times in modified gravity frameworks [146, 147, 148, 149]. In contrast, Ajmal and Sharif (2025) [150] found complete instability $(v_{s}^{2}>0)$ throughout cosmic evolution in a new agegraphic dark energy model within modified symmetric teleparallel theory, while Abdollahi Zadeh et al. (2018) [151] reported persistent instability in ghost dark energy within the DGP braneworld, regardless of interaction scenarios and also Liu et al. (2020) [152] reported the same for interacting ghost dark energy in a complex quintessence background. These comparisons underscore the distinct stability profile of EA-BADE and EA-NBADE within Agegraphic Dark Energy framework.
@@
Statefinder Parameter: The analysis of dynamical parameters in Figure 14 reveals the evolutionary trajectory in the $(r,s^{*})$ plane. During early cosmic epochs, both cosmological frameworks exhibit characteristics defined by $(r>1)$ and $(s^{*}<0)$, corresponding to quintessence behavior. The evolutionary path intersects the $\Lambda$CDM fixed point at $(r,s^{*})=(1,0)$, subsequently entering a domain characterized by $(r<1)$ and $(s^{*}>0)$, which is indicative of Chaplygin gas (Phantom-like) behavior. The dynamical evolution portrayed in Figure 15 through the $(r,q)$ plane demonstrates the cosmic transition from decelerated to accelerated expansion. This transformation reflects the shift from matter dominance to dark energy prevalence. Notable is the absence of convergence to either the steady-state configuration $(r=1,q=-1)$ or the SCDM point $(r=1,q=\frac{1}{2})$. Instead, the asymptotic behavior of both models tends toward the de Sitter state, characterized by $q=-1$. In the context of Einstein-Aether gravity reconstruction, the BADE framework exhibits more rapid convergence compared to its NBADE counterpart. These behaviors align with recent studies on statefinder diagnostics in dark energy and modified gravity models, particularly in Barrow agegraphic and Einstein-Aether frameworks, which report similar quintessence-to-phantom transitions and de Sitter convergence in the $(r,s^{*})$ and $(r,q)$ planes [50, 153, 154, 155, 156].
An important dynamical feature observed in both reconstructed Einstein-Aether gravity models—corresponding to BADE and NBADE—is the apparent transition of the effective equation of state (EoS) parameter $\omega_{DE}$ from a quintessence regime ($\omega_{DE}>-1$) at present to a phantom regime ($\omega_{DE}<-1$) in the future. While such a transition is a known characteristic of various dynamical dark energy (DDE) models [50, 54], it is crucial to determine whether this behavior in our framework exhibits any model-specific features—such as a distinct transition redshift, the sharpness or smoothness of the crossing, or dependence on model parameters—that may offer discriminative power against other scenarios. In our reconstructed EA-BADE model, the EoS parameter smoothly evolves from $\omega_{DE}\approx-0.86$ at $z=0$ to $\omega_{DE}\approx-1.05$ in the far future ($z\to-1$). For the EA-NBADE model, the present value is slightly closer to the phantom divide with $\omega_{DE}\approx-0.91$ at $z=0$, evolving to $\omega_{DE}\approx-1.12$ as $z\to-1$. Interestingly, although other diagnostic tools—such as the $(r,s)$ statefinder trajectories and the $\text{Om}(z)$ plots—suggest a late-time shift toward phantom-like behavior, the direct evolution of $\omega_{DE}(z)$ in both models does not clearly exhibit a definitive crossing of the phantom divide within the accessible redshift range. This apparent tension can be interpreted as a subtle signal of an asymptotic approach toward the phantom regime rather than a sharp transition. In particular, the value of $\omega_{DE}$ decreases smoothly and increasingly deviates from $-1$ in the negative redshift direction, but without showing a pronounced crossing point—suggesting that the models may mimic phantom behavior in diagnostics without realizing a full quintessence-to-phantom transition in the strict sense. This comparative behavior is illustrated in Figure 10(c), where we have plotted $\omega_{DE}(z)$ for EA-BADE and EA-NBADE alongside benchmark models such as CPL and Holographic Dark Energy. Notably, in parameter space, the proximity to the phantom regime is influenced by the Barrow exponent $\Delta$, which modulates the evolution rate of the agegraphic energy density. Larger $\Delta$ values tend to delay the approach to the phantom-like regime and yield a smoother evolution of $\omega_{DE}$, a feature that can potentially be tested with upcoming precision surveys like Euclid and SKA. The dependence on $\Delta$ arises uniquely from the Barrow entropy-induced modification to the agegraphic energy density, distinguishing this framework from generic quintom models that rely on multiple scalar degrees of freedom. Therefore, while the quintessence-to-phantom trend is not exclusive to our models, the smooth asymptotic evolution toward phantom behavior—combined with model-specific controls such as $\Delta$—provides a distinctive signature of the EA-BADE and EA-NBADE reconstructions, even in the absence of a strict phantom crossing in $\omega_{DE}(z)$.
@@
Comparative Analysis and Distinctive Features of EA-BADE/NBADE Models: To address the unique characteristics of our combined Einstein-Aether (EA) gravity with Barrow and non-Barrow entropic dark energy models, we present a comprehensive comparative analysis with existing modified gravity and dark energy frameworks. While several cosmological models can reproduce similar behaviors in individual diagnostic parameters, the combined EA-BADE/NBADE framework offers distinct observational signatures that differentiate it from other theoretical constructs. For instance, the quintessence-like behavior observed in our $\omega_{DE}$ analysis is common in scalar field models like quintessence [158] and k-essence [19]. However, unlike these models, our framework exhibits a distinctive evolution pattern in the statefinder plane that transitions from quintessence-like behavior through the $\Lambda$CDM fixed point to Chaplygin gas-like characteristics. This evolutionary trajectory is markedly different from that of standard scalar field models, which typically remain confined to a specific region of the statefinder plane [86].
The partial stability indicated by our $v_{s}^{2}$ analysis is particularly noteworthy when compared to other modified gravity theories. While $f(R)$ gravity models often demonstrate consistent stability across all redshift ranges [38], our EA-BADE/NBADE models exhibit a redshift-dependent stability profile with a transition at $z\approx 1$. This characteristic stability transition could serve as a distinctive observational signature for discriminating our models from other modified gravity frameworks. Furthermore, the behavior of the Om$(z)$ diagnostic reveals another unique aspect of our models. The sharp transition in the BADE model from phantom-like behavior at high redshifts to quintessence-like characteristics at low redshifts differs significantly from the behavior observed in scalar-tensor theories [159] and Galileon models [160], which typically exhibit more gradual transitions or consistent behavior across redshift ranges. This distinct feature provides a potential observational test for the EA-BADE model, as future high-precision measurements of the expansion history could detect this characteristic transition. The cosmic age estimates derived from our models ($14.00^{+0.19}_{-0.17}$ Gyr for BADE and $14.23^{+0.17}_{-0.19}$ Gyr for NBADE) align well with current observational constraints while emerging from fundamentally different physical mechanisms than standard $\Lambda$CDM cosmology. It is important to acknowledge that certain individual features observed in our models can be reproduced by other theoretical frameworks. For example:
  1. 1.
    Deceleration Parameter Evolution: The transition redshift values we obtain ($z_{t}\approx 0.7568$ for BADE and $z_{t}\approx 0.6561$ for NBADE) fall within the range predicted by various modified gravity theories, including $f(T)$ gravity [161] and Brans-Dicke theory [162]. However, the specific combination of transition redshift and present-day deceleration parameter values ($q_{0}=-0.4985$ for BADE and $q_{0}=-0.2759$ for NBADE) represents a distinctive signature of our models.
  2. 2.
    Jerk, Snap, and Lerk Parameters: The positivity of the jerk parameter across all redshifts is a feature shared with many dark energy models. However, the specific behavior of the snap parameter remaining consistently positive in both models, while the lerk parameter exhibits sign changes in the NBADE model, creates a unique cosmographic fingerprint that distinguishes our framework from other alternative gravity theories such as DGP [163] and Gauss-Bonnet gravity [164].
  3. 3.
    Equation of State Parameter: The present-day values we obtain ($\omega_{DE0}=-0.8718$ for BADE and $\omega_{DE0}=-0.7841$ for NBADE) are within the quintessence regime, similar to many dynamical dark energy models. However, the particular evolutionary trajectory of $\omega_{DE}(z)$ in our models, especially when considered alongside the $\omega^{\prime}_{DE}-\omega_{DE}$ phase plane analysis, exhibits distinctive features. Unlike standard quintessence models that typically show a thawing or freezing behavior [145], our models demonstrate a more complex evolutionary pattern in the phase plane.
  4. 4.
    Density Parameter Evolution: The matter-dark energy equality redshift of $z\approx 0.5$ is consistent with standard cosmological models. However, when analyzed in conjunction with our $Om(z)$ diagnostic, which shows model-specific deviations from $\Lambda$CDM at different redshift ranges, our framework presents a unique cosmological signature.
  5. 5.
    Statefinder Diagnostics: While individual segments of the statefinder evolutionary trajectories can be reproduced by various modified gravity and dark energy models, the complete path through the $(r,s^{*})$ and $(r,q)$ planes—beginning in the quintessence region, transitioning through the $\Lambda$CDM point, and ending in the Chaplygin gas region—represents a distinctive feature of our EA-BADE/NBADE framework. This particular evolutionary sequence is not typically observed in either pure modified gravity models like $f(R)$ theories [165] or standard dark energy models like phantom fields [166].
A key aspect that truly distinguishes our approach is the specific coupling mechanism between Einstein-Aether gravity and entropic dark energy formulations. This coupling, characterized by the parameter $M$ (measured as $M=0.771_{-0.166}^{+0.163}$ for BADE and $M=0.553_{-0.243}^{+0.270}$ for NBADE using Pantheon+SH0ES data), introduces unique effects on cosmological evolution that cannot be readily replicated by either pure modified gravity or pure dark energy models. The Barrow exponent parameter $\Delta$ in our models ($\Delta=0.324_{-0.201}^{+0.322}$ for BADE and $\Delta=0.467_{-0.286}^{+0.351}$ for NBADE using Pantheon+SH0ES data) introduces an additional layer of complexity that distinguishes our framework from conventional approaches. This parameter, arising from considerations of modified black hole entropy due to quantum-gravitational effects, introduces a fundamentally different physical mechanism for cosmic acceleration compared to standard scalar field or modified gravity approaches.
Table 7: Comparative cosmological diagnostics for dark energy models
Model $\omega_{DE}$ Behavior Ph. Cross Om$(z)$ Shape $(r,s)$ Trend Ref.
EA-BADE $-1<\omega<-0.7\to\omega<-1$ Yes ($z\approx-0.5$) Concave up ($z<1$), down ($z>1$) $(1,1)\to(1,0)$ (de Sitter)
EA-NBADE $-1<\omega<-0.8\to\omega<-1$ Yes ($z\approx-0.6$) Flat ($z<0.5$), incr. ($z>0.5$) $(1,1)\to(1,0)$ (de Sitter)
Quintessence $-1<\omega<-0.3$ No Concave down (all $z$) $(r<1,s>0)\to(1,0)$ [167]
Phantom $\omega<-1$ No Concave up (all $z$) $(r>1,s<0)\to(1,0)$ [22]
f(R) Gravity $-1.2<\omega_{\text{eff}}<-0.8$ Possible Convex/concave Smooth to $(1,0)$ [168]
Tsallis HDE $\omega\to<-1$ ($\delta$ tuned) Yes ($\delta>1$) Curved ($\delta$-dep.) Distinct ($r\neq 1,s\neq 0$) [169]
Chaplygin Gas $\omega\to-1$ (late) No Flat (late) $(1,1)\to(1,0)$ [170]
When these distinctive features are considered collectively rather than individually, our EA-BADE/NBADE framework emerges as a novel theoretical construct with unique observational signatures. Future high-precision cosmological observations, particularly those targeting the expansion history and growth of structure at intermediate redshifts ($2.00.5<z<$ 2.0), would be particularly valuable for testing the distinctive predictions of our models against other competing theoretical frameworks. In conclusion, while individual cosmological parameters in our models may exhibit behaviors similar to those in other theoretical frameworks, the specific combination of parameter values, evolutionary trajectories, and underlying physical mechanisms in the EA-BADE/NBADE models creates a distinctive cosmological signature that can be tested against future observational data. This holistic approach to model discrimination, rather than relying on individual diagnostic parameters, provides a more robust framework for evaluating the viability of our theoretical constructs against the broader landscape of cosmological models.
7 Conclusions
In this work, we performed a comprehensive reconstruction of Einstein-Aether (EA) gravity by integrating two modified agegraphic dark energy models—the Barrow Agegraphic Dark Energy and its extension, the New Barrow Agegraphic Dark Energy. The reconstruction framework involved equating the energy densities of BADE and NBADE with that of EA gravity to derive the form of the function $F(K)$, which characterizes the aether kinetic term. This approach provides a novel entropy-motivated perspective to explore dynamical dark energy in the Einstein-Aether context. We showed that the reconstructed $F(K)$ functions exhibit distinct monotonic behaviors depending on the model and parameter choices, particularly the Barrow exponent $\Delta$. While BADE produces an increasing $F(K)$ with respect to $K$, NBADE yields a decreasing profile, reflecting differing dynamical behavior rooted in the entropy assumptions.
Using a robust statistical analysis based on Markov Chain Monte Carlo (MCMC) techniques and observational constraints from CC+BAO and Pantheon+SH0ES data, we constrained the key model parameters, including $\Omega_{m0}$, $\omega_{m}$, $\Delta$, $M$, and $t_{\text{age}}$. The models produce Hubble parameter evolution in good agreement with observational measurements, though mild tensions with SH0ES remain—similar to what is observed in other non-$\Lambda$CDM models. The transition from decelerated to accelerated expansion is observed near $z_{t}\approx 0.7568$ (BADE) and $z_{t}\approx 0.6561$ (NBADE), consistent with constraints from SNIa, BAO, and CMB data. Higher-order kinematic indicators—jerk, snap, and lerk—suggest a dynamic acceleration history, and the statefinder diagnostics $(r,s^{*})$ and $(r,q)$ place the models near quintessence and Chaplygin-like regimes, asymptotically approaching de Sitter behavior.
The dark energy equation of state $\omega_{DE}$ remains negative across the evolution history, indicative of persistent acceleration. Both models exhibit a current quintessence-like behavior transitioning toward a phantom regime in the far future. This quintessence-to-phantom evolution is also captured in the $\omega_{DE}^{\prime}$$\omega_{DE}$ phase space and Om(z) diagnostics. Although this feature is common among various dynamical dark energy models, the redshift of transition and the corresponding phase space trajectories in our EA-BADE and EA-NBADE frameworks may provide distinguishing signatures when evaluated against future high-precision datasets. Stability analysis using the squared sound speed $v_{s}^{2}$ reveals partial epoch-dependent stability, with the models being stable at higher redshifts and exhibiting instabilities near the current epoch. This behavior is consistent with some other noncanonical DE models and signals a need for further theoretical scrutiny regarding perturbative behavior in modified gravity scenarios.
In comparative context, while many of the observed features—such as late-time acceleration, freezing behavior in the $\omega_{DE}^{\prime}$$\omega_{DE}$ plane, and redshift-dependent Om diagnostics—are shared by other well-studied models (e.g., CPL, Holographic, or quintom-type models), the EA-BADE and EA-NBADE reconstructions offer a distinct pathway by incorporating entropy-based corrections to agegraphic energy density. The Barrow exponent $\Delta$, in particular, plays a critical role in modulating model behavior and could serve as a key discriminator in future observational tests.
Although the models do not yet demonstrate a decisive improvement over $\Lambda$CDM in terms of statistical significance, their ability to encode entropy deformation in a covariant gravitational framework provides a fertile ground for further theoretical developments. The models serve as useful phenomenological laboratories to probe the interface of modified gravity, nonextensive thermodynamics, and cosmic acceleration. Future research should examine possible time variation of the Barrow exponent $\Delta$—as suggested by recent studies—and explore generalized entropic frameworks such as the Nojiri-Odintsov-Paul entropy formalism. Additionally, incorporating interaction terms between dark energy and dark matter, as well as leveraging more extensive datasets (e.g., CMB, LSS, and gravitational wave backgrounds), could sharpen parameter constraints and assess the long-term viability of the models. Finally, implementing these models in large-scale structure simulations would allow for direct comparison with $\Lambda$CDM predictions and further illuminate the potential of Einstein-Aether gravity in the context of cosmic evolution. In summary, while not conclusively superior to standard cosmological models, the EA-BADE and EA-NBADE frameworks represent a promising and theoretically motivated extension of Einstein-Aether gravity. Their compatibility with observational data and capacity to encapsulate non-standard thermodynamic corrections position them as viable candidates for further exploration in the quest to understand the nature of dark energy and the fundamental structure of spacetime.